# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ Abstract
The transition from 5G networks to 6G highlights a significant demand for machine learning (ML). Deep learning models, in particular, have seen wide application in mobile networking and communications to support advanced services in emerging wireless environments, such as smart healthcare, smart grids, autonomous vehicles, aerial platforms, digital twins, and the metaverse. The rapid expansion of Internet-of-Things (IoT) devices, many with limited computational capabilities, has accelerated the development of tiny machine learning (TinyML) and resource-efficient ML approaches for cost-effective services. However, the deployment of large-scale machine learning (LargeML) solutions require major computing resources and complex management strategies to support extensive IoT services and ML-generated content applications. Consequently, the integration of TinyML and LargeML is projected as a promising approach for future seamless connectivity and efficient resource management.
Although the integration of TinyML and LargeML shows abundant potential, several challenges persist, including performance optimization, practical deployment strategies, effective resource management, and security considerations. In this survey, we review and analyze the latest research aimed at enabling the integration of TinyML and LargeML models for the realization of smart services and applications in future 6G networks and beyond. The paper concludes by outlining critical challenges and identifying future research directions for the holistic integration of TinyML and LargeML in next-generation wireless networks.
@@
6G, artificial intelligence (AI), deep learning (DL), distributed learning, federated learning (FL), Internet-of-Things (IoT), large machine learning (LargeML), machine learning (ML), tiny machine learning (TinyML).
@@ I Introduction
The evolution of communication technologies has been characterized by extraordinary advancements every decade. From the first-generation (1G) analog systems, which introduced mobile voice communications, to fifth-generation (5G) networks currently transforming industries with ultra-fast speeds and low latency, each generation has delivered major transformative impacts [1]. Looking ahead, sixth-generation (6G) networks are expected to push technological boundaries further, offering unprecedented capabilities and driving global innovation. The core features of 6G will clearly differentiate it from previous generations (Fig. 1), positioning the paradigm as a transformative technology for diverse applications. These features address the escalating demands for higher data rates, superior reliability, efficient spectrum management, and superintelligence capabilities. Furthermore, 6G aims to integrate advanced artificial intelligence (AI) and ML technologies seamlessly, expanding the possibilities of wireless networks and significantly improving the delivery of high-end applications and services to users.
Figure 1: Transition from 5G to 6G and beyond (B6G): From target throughputs to use cases, core technologies, and emerging applications.
The integration of AI and ML into 6G will enable new levels of automation and intelligence, unlocking numerous new applications and high-end services [2]. AI already operates on multiple network layers to improve dynamic resource management, predict maintenance needs, and personalize user experiences. Advances in ML, particularly deep learning (DL) architectures that execute computer vision and natural language processing (NLP) tasks [3], are expected to support high-fidelity holographic communications. This capability will allow users to interact in immersive three-dimensional environments and could revolutionize remote collaboration. In a hyper-connected future, 6G will unlock the full potential of Internet-of-Things (IoT) and mobile big data to drive innovation and lead the world into smarter, deeper connectivity.
@@ I-A Context and Motivation
In the last decade, AI has converged around two contrasting approaches: TinyML and Large ML (LargeML). TinyML focuses on creating efficient models for resource-constrained edge devices, allowing real-time analytics close to data sources (i.e., bringing analytics closer to data sources) [4]. In contrast, LargeML leverages powerful cloud servers to handle complex, data-intensive tasks, exemplified by models like GPT-4 and BERT, which excel in applications such as sentiment analysis and creative text generation [5] (e.g., Large Language Models (LLMs), generative AI (GenAI), and Agentic AI).
However, LargeML’s reliance on significant computational resources limits its real-time processing abilities on constrained devices and raises privacy concerns. TinyML mitigates these issues by pre-processing data at the edge, while LargeML can deploy compressed models on edge devices through knowledge distillation [6]. This collaborative learning strategy permits more flexible deployment and operation of IoT applications and services. TinyML benefits from the sophisticated learning capabilities of LargeML to improve accuracy, whereas LargeML gains from the real-time processing potential and privacy-preserving features of TinyML. Table I presents a concise comparison of TinyML and LargeML.
The integration of TinyML and LargeML addresses these individual limitations by combining their strengths, enhancing AI capabilities in 6G and future networks. For instance, wearable health monitors can use TinyML for real-time vital sign analysis [7], while LargeML can detect medical risks using historical data. Similarly, in virtual environments, TinyML can process wearable Electroencephalography (EEG) data to gauge emotional states [8], with LargeML translating this into actions within the metaverse [9].
Despite these advantages, the integration faces challenges, including increased security risks from numerous connected devices and ethical concerns regarding AI explainability and data privacy. In summary, leveraging the unique strengths of TinyML and LargeML can lead to seamless device-to-server interactions and drive advancements in data-driven innovation and intelligent living environments.
@@ I-B State-of-the-Art and Contributions
The AI ecosystem is evolving rapidly, driven in part by two distinct yet complementary paradigms: TinyML and LargeML. TinyML enables resource-limited edge devices to perform on-device ML tasks with local data, whereas LargeML employs powerful server infrastructures to train complex models on vast datasets. As the IoT continues to evolve and the advent of 6G approaches, the collaborative potential of these two different learning strategies presents a viable path on which to increase the performance of IoT applications and 6G network operations, especially under challenging conditions. This section reviews the recent survey literature on TinyML and LargeML systems, highlighting their contributions and limitations in the context of 6G and its envisioned applications.
Over the past decade, numerous survey works have examined the role of TinyML in edge computing. Key contributions [10, 11, 12, 13, 14, 15] provide comprehensive overviews of TinyML, covering its fundamentals, development tools, applications, and future directions. These surveys highlight both the transformative potential and widespread use of TinyML in domains such as industrial IoT, smart healthcare, autonomous driving, environmental monitoring, public safety, human-machine interaction, agriculture, and emergency response. A significant contribution noted in [10] is the concept of hardware-software cooperative design for efficient TinyML implementation. This survey highlights the necessity of optimizing both hardware and software components concurrently for efficient TinyML execution on resource-constrained devices. Coordinated design enhances model efficiency while minimizing power consumption and computational overhead. Similarly, the survey in [12] offers a systematic overview of existing TinyML development tools, including hardware platforms, software frameworks, and supporting libraries. These resources assist AI developers and specialists in designing and deploying TinyML models at the network edge. Other surveys [11, 14, 15] explore a range of TinyML applications across domains such as anomaly detection, predictive maintenance, and intelligent sensor networks. For example, anomaly detection allows TinyML models to continuously monitor sensor data and identify deviations from normal operating behavior, supporting timely maintenance and minimizing unplanned downtime. In predictive maintenance, TinyML models forecast potential equipment failures, improving resource allocation and operational efficiency through proactive maintenance strategies. Intelligent sensor networks use TinyML to process data locally at the edge, enabling quicker decision-making and near real-time responsiveness. These applications illustrate the broad potential of TinyMl to transform sectors such as industrial automation, environmental monitoring, and smart infrastructure.
@@
[10] Highlights hardware-software co-design for optimizing TinyML frameworks on resource-limited devices and reviews ML algorithms optimized for low-power deployment. Concentrates on TinyML and 5G, without discussion of integration with LargeML in 6G systems.
[11] Introduces a classification system for TinyML applications based on an extensive literature review, exploring benefits and use cases. Does not address TinyML–LargeML integration or the specific challenges of 6G network integration.
[12] Surveys TinyML frameworks, development tools, enablers, and applications and outlines current research challenges. Lacks discussion of TinyML–LargeML integration in the context of 6G networks.
@@
[15] Reviews TinyML research, workflows, algorithms, and development tools, with an emphasis on recent ML advancements. Lacks analysis of integration with LargeML in the context of 6G networks.
@@
[17] Examines LLM evaluation methods, benchmarks, and performance across domains. Omits discussion of TinyML–LargeML integration and analysis of challenges in enhancing 6G applications.
[18] Analyzes fine-tuning algorithms for LargeML, detailing performance metrics and deployment scenarios. Does not investigate the integration of fine-tuning algorithms with TinyML for 6G applications.
@@
[20] Outlines architectures, techniques, and challenges in pre-trained language models (PLMs). Concentrates on natural language processing (NLP) applications of PLMs; lacks discussion of TinyML integration in 6G systems.
@@
This work Offers the first survey of the integration of TinyML and LargeML for 6G systems. Specifically: Introduces fundamentals of TinyML and LargeML. Discusses motivations and requirements for integrating TinyML and LargeML in 6G. Proposes bidirectional integration frameworks for enhancing future network performance and capabilities. Explores potential applications of integrated TinyML–LargeML systems. Identifies challenges and future research directions for TinyML–LargeML integration.
In contrast, many studies on LargeML outline distinct contributions and limitations. Surveys such as [16, 17, 18, 19, 20], explore the architectures, applications, evaluation methods, and research challenges associated with LargeML. These works offer a clear view of the current state of the LargeML field. For example, the work in [16] presents a detailed analysis of transformer-based architectures, from foundational models to advanced, with a focus on LLMs. It traces the development of of LLMs, discussing training methods, application domains, and real-world deployments. Similarly, the survey in [20] explores the impact of large pre-trained language models (PLMs) on a range of NLP tasks. These surveys highlight the strong performance of LargeML in complex tasks such as text generation, machine translation, and sentiment analysis. Several reviews [17, 18, 19] examine key challenges associated with LargeML. These include explainability, where understanding model decision-making processes remains critical for real-world deployment. Notably, [18] emphasizes the need for parameter-efficient fine-tuning (PEFT) techniques to reduce computational overhead, facilitating deployment in resource-constrained environments. Addressing these challenges is essential for the responsible and effective use of LargeML across diverse applications. However, as with the surveys on TinyML, current LargeML reviews exhibit limitations when considered in the context of 6G applications. Although the computational demands of LargeML are acknowledged, the specific challenges of deployment and and coordination with LargeML within 6G networks are not examined. Given 6G’s focus on real-time processing and edge intelligence, it becomes necessary to investigate how LargeML models can be compressed or adapted for efficient operation at the network edge, potentially in conjunction with TinyML systems. Furthermore, the surveys overlook the impact of 6G-specific features such as novel communication protocols and strict latency constraints on the training and deployment of LargeML models.
The preceding discussion highlights a significant gap in comprehensive surveys that examine the integration of TinyML and LargeML in the context of 6G and beyond. Although recent progress in LargeML integration [21, 22] is critical for developing efficient and intelligent 6G networks, these advancements have not been thoroughly examined. Existing studies focus on isolated aspects of either TinyML or LargeML [23], failing to provide a cohesive analysis of their combined potential to address emerging network challenges. The present survey aims to fill this gap by presenting a comprehensive survey on the integration of TinyML and LargeML for 6G and future network paradigms. A key contribution is a detailed examination of the motivations and requirements behind such integration. We begin by reviewing the foundations and recent advancements in both TinyML and LargeML, highlighting their respective strengths. We then propose efficient bidirectional integration strategies and asses their relevance to future network architectures. In addition, we examine the real-world applications of integrated TinyML–LargeML systems in areas such as smart healthcare, autonomous vehicles, and industrial IoT. These case studies illustrate the potential to improve communication efficiency, optimize resource use, and scale systems systems effectively. Finally, we identify major unresolved challenges in the literature and propose research directions to guide future work. The features and contributions of this work are summarized below:
@@
  • Background: This work first provides the background of TinyML and LargeML, accordingly highlighting their individual advancements and potential cooperation.
@@
  • Efficient solutions: We propose and discuss efficient bidirectional integration solutions, aiming to enhance the performance and capabilities of future networks.
@@
  • Discussion of challenges and future research: From our extensive review, we identify critical challenges and propose several potential directions for future research, including resource allocation, communication efficiency, interplay with LargeML, and standardizations.
I-C Paper Organization
The remainder of the paper is structured as follows. Section II provides an overview of TinyML and LargeML, outlining their key concepts and recent developments Section III examines the motivations and requirements for integrating TinyML and LargeML in the context of 6G networks. Section IV presents a comprehensive review of efficient, bi-directional integration approaches for 6G and beyond, such as transfer learning (TL), federated transfer learning (FTL), split learning (SL), and federated split learning (FSL). Section V explores the applications of integrated TinyML and LargeML systems in domains such as data privacy and network security, network management, zero-touch networks, and the brain-level metaverse. In Section VI, we summarize insights from earlier sections and discuss existing challenges and potential directions for future research before ending with Section VII.
@@ II Overview of TinyML and LargeML
Figure 2: Overview of TinyML and LargeML: life cycle, design, operation, and performance evaluation.
II-A TinyML
TinyML refers to a lightweight, resource-constrained class of ML designed for environments with limited memory and processing capabilities [4]. Despite its compact form, TinyML performs effectively in a wide range of communication tasks; Table III details its characteristics.
@@ II-A TinyML
Model size: (1) Compression techniques (pruning, quantization, knowledge distillation), (2) Memory and storage design (kilobytes to megabytes), (3) Operational modes (centralized, distributed, decentralized). Architecture: (1) Convolutional neural network (CNN) for image processing, (2) Recurrent neural network (RNN) for sequential data, (3) Fully connected networks for certain tasks, (4) Transformer [24] for data generation, context awareness, translation, and summarization. Training data: (1) Quality assurance and data labeling, (2) Data augmentation techniques (rotation, scaling, flipping), (3) Fine-tuning pre-trained models on large datasets. Algorithms: Ten types [11]: supervised, weakly supervised, unsupervised, self-supervised, meta-learning, continual learning, deep reinforcement learning (DRL), TL, SL, and federated learning (FL). Central processing unit (CPU): Arm Cortex M family offers ultra-low-power operation but limited support for parallel processing. Graphics processing unit (GPU): Supports parallel computation and optimized sequential processing of large datasets (e.g. NVIDIA Jetson family, AMD, Intel, Arm), but is not cost-effective. Field-programmable gate array (FPGA): Enables model customization and hardware-level acceleration for complex operations and large dataset access, but lacks broad library support. Tensor processing unit (TPU): Accelerates ML workloads, particularly those involving NNs, with two representations of Edge TPUs and Cloud TPUs. Atmospheric: Low-power sensors to monitor air quality in urban, forest, and industrial areas, detecting pollutants by measuring particulate matter, carbon dioxide levels, and others. Hydrosphere: Sensors in aquatic environments to detect pollutants, measure pH, and monitor temperature changes. Biosphere: Device used for wildlife conservation, monitoring biological activity, or classifying sounds (e.g., animal habitats and anti-poaching surveillance). Lithosphere: Applied in agriculture for environmental monitoring (e.g., soil moisture, nutrient levels, pests) or earthquake detection. Human-related: Wearables for health monitoring (e.g., blood pressure) and assistive technologies for humans (e.g., mobility). Mobile device: TensorFlow Lite for model optimization and PyTorch for programming. Embedded devices: TensorFlow Lite for model optimization, PyTorch for programming, and Qeexo AutoML for end-to-end customization. Edge devices: PyTorch for programming, NanoEdgeAIStudio for low/no-code solution, and Imagimob/Edge Impulse for end-to-end customization. Microcontroller: Arm NN for network architecture optimization, MinUn for bandwidth/memory management, STM32CubeA and uTensor for pre-trained model conversion, and NXP eIQ as a development environment.
According to [25], the TinyML life cycle consists of three stages:
  • Stage 1: Data collection from peripheral devices, such as sensors or production environments (real or simulated).
  • Stage 2: Model deployment using quantization and compression techniques to address hardware heterogeneity and resource constraints.
  • Stage 3: Model operationalization by converting the trained model into a generic, interpretable form suitable for target devices.
TinyML relies on application-specific data formats, prioritizing model customization for limited device capacity and efficient data transmission. The next sections detail these design strategies and operational considerations (Fig. 2 for short overview).

### v2

@@ Abstract
The evolution from fifth-generation (5G) to sixth-generation (6G) networks is driving an unprecedented demand for advanced machine learning (ML) solutions. Deep learning has already demonstrated significant impact across mobile networking and communication systems, enabling intelligent services such as smart healthcare, smart grids, autonomous vehicles, aerial platforms, digital twins, and the metaverse. At the same time, the rapid proliferation of resource‑constrained Internet‑of‑Things (IoT) devices has accelerated the adoption of tiny machine learning (TinyML) for efficient on‑device intelligence, while large machine learning (LargeML) models continue to require substantial computational resources to support large‑scale IoT services and ML‑generated content. These trends highlight the need for a unified framework that integrates TinyML and LargeML to achieve seamless connectivity, scalable intelligence, and efficient resource management in future 6G systems. This survey provides a comprehensive review of recent advances enabling the integration of TinyML and LargeML in next‑generation wireless networks. In particular, we (i) provide an overview of TinyML and LargeML, (ii) analyze the motivations and requirements for unifying these paradigms within the 6G context, (iii) examine efficient bidirectional integration approaches, (iv) review state‑of‑the‑art solutions and their applicability to emerging 6G services, and (v) identify key challenges related to performance optimization, deployment feasibility, resource orchestration, and security. Finally, we outline promising research directions to guide the holistic integration of TinyML and LargeML for intelligent, scalable, and energy‑efficient 6G networks and beyond.
@@
6G, artificial intelligence (AI), deep learning (DL), federated learning (FL), Internet-of-Things (IoT), large machine learning (LargeML), tiny machine learning (TinyML).
@@ I Introduction
The evolution of communication technologies has been characterized by extraordinary advances every decade. From the first-generation (1G) analog systems, which introduced mobile voice communications, to fifth-generation (5G) networks currently transforming industries with ultra-fast speeds and low latency, each generation has delivered major transformative impacts [35]. Looking ahead, sixth-generation (6G) networks are expected to push technological boundaries further, offering unprecedented capabilities and driving global innovation. The core features of 6G will clearly differentiate it from previous generations (Fig. 1), positioning the paradigm as a transformative technology for diverse applications. These features address the escalating demands for higher data rates, superior reliability, efficient spectrum management, and superintelligence capabilities. Furthermore, 6G aims to seamlessly integrate advanced artificial intelligence (AI) and machine learning (ML) technologies, expanding the possibilities of wireless networks and significantly improving the delivery of high-end applications and services to users.
The integration of AI and ML into 6G will enable new levels of automation and intelligence, unlocking numerous new applications and high-end services [205]. AI already operates on multiple network layers to improve dynamic resource management, predict maintenance needs, and personalize user experiences. Advances in ML, particularly deep learning (DL) architectures that execute computer vision and natural language processing (NLP) tasks, are expected to support high-fidelity holographic communications. This capability will allow users to interact in immersive three-dimensional environments and could revolutionize remote collaboration. In a hyper-connected future, 6G will unlock the full potential of the Internet-of-Things (IoT) and mobile big data to drive innovation and lead the world into smarter, deeper connectivity.
@@ I-A Context and Motivation
Figure 1: Transition from 5G to 6G and beyond 6G (B6G): Target throughput, use cases, core technologies, and applications.
@@
In the last decade, AI has converged around two contrasting approaches: tiny machine learning (TinyML) and large-scale machine learning (LargeML). TinyML focuses on creating efficient models for resource-constrained edge devices, allowing real-time analytics close to data sources (i.e., bringing analytics closer to data sources) [97]. In contrast, LargeML leverages powerful cloud servers to handle complex, data-intensive tasks, exemplified by models like GPT-4 and BERT, which excel in applications such as sentiment analysis and creative text generation [193] (e.g., Large Language Models (LLMs), generative AI (GenAI), and Agentic AI).
However, LargeML’s heavy computational demands limit real‑time use on constrained devices and raise privacy concerns, while TinyML mitigates these issues by pre‑processing data at the edge and enabling LargeML to deploy compressed models through knowledge distillation [213]. This collaborative strategy improves flexibility for IoT services: TinyML gains accuracy from LargeML’s advanced learning, and LargeML benefits from TinyML’s real‑time processing and privacy preservation, as summarized in Table I. Their integration strengthens AI capabilities in 6G systems, for example, wearable health monitors can analyze vital signs locally with TinyML [167] while LargeML detects medical risks from historical data, and TinyML can process Electroencephalography (EEG) signals for emotional state recognition [121] while LargeML translates them into metaverse actions [183].
Despite these advantages, challenges remain, including heightened security risks from massive device connectivity and ethical concerns around explainability and data privacy, yet combining TinyML and LargeML ultimately enables seamless device-server interaction and advances in data-driven innovation and intelligent living environments.
@@ I-B State-of-the-Art and Contributions
Over the past decade, numerous surveys have examined TinyML’s role in edge computing, offering comprehensive overviews of its fundamentals, development tools, applications, and future directions [42, 1, 137, 136, 168, 23], with key contributions including hardware–software co‑design for efficient deployment on constrained devices [42] and systematic reviews of TinyML toolchains spanning hardware platforms, software frameworks, and libraries [137]. Other works [1, 168, 23] highlight TinyML’s broad application landscape, from anomaly detection and predictive maintenance to intelligent sensor networks, where models monitor sensor data for deviations, forecast equipment failures, and enable fast, local decision‑making. Collectively, these surveys highlight TinyML’s transformative potential across industrial automation, environmental monitoring, smart infrastructure, and other data‑driven domains.
@@
In contrast, LargeML surveys outline complementary strengths and limitations, with works such as [135, 24, 56, 207, 115] examining architectures, applications, evaluation methods, and open challenges. Reviews like [135] analyze transformer-based models and LLM development, while [115] highlights the impact of large pre-trained language models (PLMs) on diverse NLP tasks, emphasizing their strong performance in text generation, translation, and sentiment analysis. Other surveys [24, 56, 207] focus on challenges such as explainability and the need for parameter-efficient fine-tuning (PEFT) [56] to reduce computational overhead. However, despite their breadth, current LargeML surveys do not address 6G‑specific constraints: they overlook deployment and coordination challenges in real‑time, latency‑sensitive, edge‑intelligent networks, and the need to compress or adapt LargeML models for efficient operation alongside TinyML systems. They also omit the implications of emerging 6G communication protocols and strict latency requirements for training and inference. Table II summarizes these surveys and situates the contributions of this work.
@@
[42] Highlights hardware-software co-design for optimizing TinyML frameworks on resource-limited devices and reviews ML algorithms optimized for low-power deployment. Focuses on TinyML and 5G, without discussion of integration with LargeML in 6G systems.
[1] Introduces a classification system for TinyML applications based on an extensive literature review, exploring benefits and use cases. Does not address TinyML–LargeML integration or the specific challenges of 6G integration.
[137] Surveys TinyML frameworks, development tools, enablers, and applications, and outlines current research challenges. Lacks discussion of TinyML–LargeML integration in the context of 6G networks.
@@
[23] Reviews TinyML research, workflows, algorithms, and development tools, with an emphasis on recent ML advances. Lacks analysis of integration with LargeML in the context of 6G networks.
@@
[24] Examines LLM evaluation methods, benchmarks, and performance across domains. Omits discussion of ML integration and analysis of challenges in 6G applications.
[56] Analyzes fine-tuning algorithms for LargeML, detailing performance metrics and deployment scenarios. Lacks the investigation of fine-tuning algorithms with TinyML for 6G applications.
@@
[115] Outlines architectures, techniques, and challenges in PLMs. Focuses on NLP applications of PLMs; lacks discussion of TinyML integration in 6G systems.
@@
This work Offers the first survey of the integration of TinyML and LargeML for 6G systems. Specifically, (i) Introduces fundamentals of TinyML and LargeML; (ii) Discusses motivations and requirements for integrating TinyML and LargeML in 6G; (iii) Proposes bidirectional integration frameworks for enhancing future network performance and capabilities; (iv) Explores potential applications of integrated TinyML–LargeML systems; and (v) Identifies challenges and future research directions for TinyML–LargeML integration.
The previous discussion reveals a significant gap in surveys addressing the integration of TinyML and LargeML in the context of 6G and beyond. Although recent advancements in LargeML [214, 139] are vital for developing efficient 6G networks, they have not been thoroughly analyzed. Most studies focus on either TinyML or LargeML [179], lacking a cohesive examination of their combined potential to tackle emerging network challenges. This survey aims to fill that gap by exploring the integration of TinyML and LargeML for 6G and future networks. A key contribution of this work is a thorough exploration of the motivations and requirements for such integration. Our contributions are summarized below:
  1. 1.
@@
    Background: This work first provides the background of TinyML and LargeML, highlighting their individual advances and potential for cooperation.
  2. 2.
@@
  3. 3.
    Efficient solutions: We propose and discuss efficient bidirectional integration solutions aimed at enhancing the performance and capabilities of future networks.
  4. 4.
@@
  5. 5.
    Discussion of challenges and future research: From our extensive review, we identify critical challenges and propose several potential directions for future research, including resource allocation, communication efficiency, interplay with LargeML, and standardization.
@@ II Overview of TinyML and LargeML
II-A Overview of TinyML
TinyML is a lightweight, resource-constrained class of ML designed for environments with limited memory and processing power [97]. TinyML supports diverse communication tasks, and its characteristics are summarized in Table III.
@@ II-A Overview of TinyML
Model size: (1) Compression methods (pruning, quantization, knowledge distillation), (2) Memory/storage design (kilobytes to megabytes), (3) Working modes (centralized, distributed, decentralized). Architecture: (1) Convolutional neural network (CNN) for image processing, (2) Recurrent neural network (RNN) for sequential data, (3) Fully connected networks for certain tasks, (4) Transformer [173] for data generation, context awareness, translation, and summarization. Training data: (1) Quality assurance and data labeling, (2) Data augmentation (rotation, scaling, flipping), and (3) Fine-tuning pre-trained models on large datasets. Algorithms: Supervised, weakly supervised, unsupervised, self-supervised, meta-learning, continual learning, deep reinforcement learning (DRL), transfer learning (TL), split learning (SL), and federated learning (FL) [1] (readers may refer to [118, 17, 185, 178, 197] for additional foundations and core concepts). Central processing unit (CPU): Arm Cortex M family offers ultra-low-power operation but limits to parallel processing. Graphics processing unit (GPU): Supports parallel computing and optimized sequential processing of large datasets (e.g., NVIDIA Jetson family, AMD, Intel, Arm), but is not cost-effective. Field-programmable gate array (FPGA): Enables model customization and hardware-level acceleration for complex operations and large dataset access, but lacks broad library support. Tensor processing unit (TPU): Accelerates ML workloads, particularly those involving NNs, with two representations of Edge TPUs and Cloud TPUs. Atmospheric: Low-power sensors to monitor air quality in urban, forest, and industrial areas, detecting pollutants by measuring particulate matter, carbon dioxide levels, and others. Hydrosphere: Sensors in aquatic environments to detect pollutants, measure pH, and monitor temperature changes. Biosphere: Device used for wildlife conservation, monitoring biological activity, or classifying sounds (e.g., animal habitats and anti-poaching surveillance). Lithosphere: Applied in agriculture for environmental monitoring (e.g., soil moisture, nutrient levels, pests) or earthquake detection. Human-related: Wearable health monitor (e.g., blood pressure) and assistive technologies for humans (e.g., mobility). Mobile device: TensorFlow Lite for model optimization and PyTorch for programming. Embedded devices: TensorFlow Lite for model optimization, PyTorch for programming, and Qeexo AutoML for end-to-end customization. Edge devices: PyTorch for programming, NanoEdgeAIStudio for low/no-code solution, and Imagimob/Edge Impulse for end-to-end customization. Microcontroller: Arm NN for network architecture optimization, MinUn for bandwidth/memory management, STM32CubeA and uTensor for pre-trained model conversion, and NXP eIQ as a development environment.
Figure 2: Overview of TinyML and LargeML, including their life cycle, design, operation, and performance evaluation.
As [189], the TinyML life cycle has three stages: (i) Stage 1: Data collection from peripheral devices (sensors or production environments); (ii) Stage 2: Model deployment using quantization and compression to address hardware heterogeneity and resource constraints; and (iii) Stage 3: Model operationalization, converting trained models into interpretable forms for target devices. TinyML relies on application-specific data formats, prioritizing model customization for device capacity and efficient data transmission (see Fig. 2 for a short overview).